Actions
Ops/Comm Scheduling meeting¶
Location : CR12 in R68 @ 15:30 pm on September 13th
Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402
Hall work¶
- DC report
- Water leak status
- Decay solenoid leak status
- Magnet test plans to Saturday.
- Hydrogen system work
Operations¶
- MOM report
- Expert led run
- http://micewww.pp.rl.ac.uk/projects/operations/wiki/_StepIV_Shutdown_Summer_
- Can we start taking data next Monday?
Tentative weekly plan¶
- Tuesday morning
- Magnets
- Tuesday afternoon (subject to waterflow)
- YK trigger fiddling (with beam if possible)
- Tuesday evening
- Magnets
- Wednesday (subject to waterflow)
- neutron monitor checks
- expert run
- beam bump tuning (Wednesday pm)
- Thursday - Sunday
- magnets
- Monday
- Begin data taking
Minutes¶
- Present : SB, CW, VB, AB, AM, VB, PF, JB, RB, CR, CM, HN
- DC report
- Done
- Water leak status
- Water leaks have been repaired but the pumps are not on since work is ongoing in the trench
- Roof system is fine. We can either run the magnet on local control (time and no temperature regulation) or complete the work
- Will work on this tomorrow when Henry is here.
- Decay solenoid leak status
- Leak has not been found. Problem could be with the air end (which has just been replaced). The leadtime for a new air end is 6-8 weeks + installation
time (Argh!). - Company is investigating if this is the problem.
- In the meantime, a different expert is coming tomorrow to check the problem. More info by the end of the week.
- Leak has not been found. Problem could be with the air end (which has just been replaced). The leadtime for a new air end is 6-8 weeks + installation
- Magnet test plans to Saturday.
- Tuesday : SSD ECE to 3T
- Wednesday : Then bring the FC in and look at M1 and SSU to tune the quench protection. At that point we need to decide whether we want to soak test to 6.1.4 settings or push the envelope a bit more.
- Hydrogen system work
- Mark Tucker needs to test some of the relief valves in the Hall between now and LH2 installation.
- Sooner rather than later
- Could be done between the user cycles or in the Wednesday maintenance slots
- R9 : Andy/wants to fit temperature sensors to FC magnet
- R9 : Target spare rack needs a good clean. Colin wants to move it to R9 which is a few week job
- Decision is needed in the next few weeks.
Operations¶
- MOM report
- Expert led run
- Hardware issues so ELR has not happened in the control room. Tests on DAQ and software was done in the remote CR.
- Workflow has been signed off using pulser triggers
- Run Control doesn't work yet
- Beamline not tested yet
- Beam bump tuning will happen on Wednesday afternoon
- http://micewww.pp.rl.ac.uk/projects/operations/wiki/_StepIV_Shutdown_Summer_
- Can we start taking data next Monday?
Updated by Boyd, Steven about 7 years ago ยท 9 revisions